OpenStudy (anonymous):

Please help me! Write the expression as the sine, cosine, or tangent of an angle. sin 9x cos x - cos 9x sin x

OpenStudy (anonymous):

sin(a - b) = sin(a)cos(b) - cos(a)sin(b) Here a = 9x and b = x

OpenStudy (anonymous):

since sin u cos v -cos u sin v=sin (u-v) =sin(9x-x)=sin(8x)

OpenStudy (anonymous):

The equation I gave is an identity for sin and cos and is good to memorize eventually as it is prevalent in math. Now, just make the substitutions and you'll be able to derive your answer.